is selected at random from the population represented by the graph. Find the probability that the member selected at random is from the shaded area of the graph. Assume the variable x is SAT Cica Reading Scaes Q 200 4x< 400 514 σ-122 200 400 800 Score The probability that the member selected at random is from the shaded area of the graph is(Round to four decimal places as needed)
Answer #1

x-μ 200-514 122 400-514 122 when x= 200, z- 2 57 when x = 400: z = x=400; z = 400-5 -0.93 p(200x<400) = p(-2.57 < z <-0.93) 0